How to solve a linear equation 2x-y=3 & 3x+y=2
Answer by richwmiller(17219)

2x-y=3
3x+y=2
add them together
5x=5
x=1
can you find y now?
plug in 1 for x in the second equation
3*1+y=2
3+y=2
y=-1